OpenStudy (anonymous):

A rectangular prism is truncated removing 30% of its height. The original height of the prism was 100 cm and it has a length of 6 cm and a width of 12 cm. What is the volume of the truncated prism?

OpenStudy (schrodinger):

The volume of a rectangular prism is length*width*height, correct? If so, V = l*w*h. V(original) = .06*.12*1.0 V(new) should be .06*.12*(?). (?) = .7(1.0) (because a 30 percent decrease will be 70% of the previous value, i.e. that value times .7) V(new) = .06*.12*.70 = ??? There's your equation. Got it now?

OpenStudy (anonymous):

Yeah thanks!

OpenStudy (schrodinger):

No problem dude.

OpenStudy (anonymous):

i don't get it @schrodinger